mar buys a bag of cookies that contains 8 chocolate chip cookies, 8 peanut butter cookies, 4 sugar cookies and 4 oatmeal cookies. What is the probability that Omar reaches in the bag and randomly selects an oatmeal cookie from the bag, eats it, then reaches back in the bag and randomly selects a sugar cookie? Give your answer as a fraction, or accurate to at least 4 decimal places. LicensePoints possible: 3 This is attempt 1 of 3.
Answer #1

Total number of cookies = 8 + 8 + 4 + 4 = 24

So,

P(Omar draws an oatmeal cookie and then a sugar cookie)

= (4 / 24) * (4 / 23)

= 2/69